# Products in spinc-cobordism

### v1

@@ 3. The $Spin^{c}$ bordism ring in low degrees
The image of the map $\Omega^{Spin}_{*}\longrightarrow\mathfrak{N}_{*}$ agrees, in degrees $\leq 31$, with subring of $MO_{*}$ generated by the elements
@@
Each of the elements $T_{i}$ defined in Proposition 3.2 is a linear combination of monomials in $\mathfrak{N}_{*}$. Those monomials are generally not individually members of $\Omega^{Spin^{c}}_{*}$: for example, $Y_{14}Y_{5}^{2}\in\mathfrak{N}_{24}$ does not lift to an element of $\Omega^{Spin^{c}}_{24}$, even though a linear combination of $Y_{14}Y_{5}^{2}$ with other monomials in degree $24$ does lift to the element $T_{24}\in\Omega^{Spin^{c}}_{24}$.
However, $Y_{i}^{2}\in\mathfrak{N}_{2i}$ lifts to the element $Z_{2i}\in\Omega^{Spin^{c}}_{2i}$, and consequently the squares of each of the monomials in each of the elements $T_{i}$ lift to $\Omega^{Spin^{c}}_{*}$. For $i=24,29,31,32$, and $33$, let $U_{2i}$ denote the element of $\Omega^{Spin^{c}}_{*}$ obtained by taking the definition of $T_{i}$ in Proposition 3.1 and replacing each instance of $Y_{n}$ with $Z_{2n}$. For example, (8) yields that
@@
The ring $\left(\Omega^{Spin^{c}}_{*}/T\right)\otimes_{\mathbb{Z}}\mathbb{F}_{2}$ was calculated by Stong [23, Proposition 11]: it is a polynomial $\mathbb{F}_{2}[\beta]$-algebra on generators in degrees $4,8,12,16,\dots$. Since $\tilde{T}$ is an ideal in $\Omega^{Spin^{c}}_{*}$, to calculate the product in the ring $\Omega^{Spin^{c}}_{*}\otimes_{\mathbb{Z}}\mathbb{F}_{2}$, it suffices to calculate
@@
Both of these types of products land in $\tilde{T}$. Since $\tilde{T}$ maps injectively under the map $\Omega^{Spin^{c}}_{*}\rightarrow\mathfrak{N}_{*}$, we can embed $\Omega^{Spin^{c}}_{*}/(2,\beta)$ into $\mathfrak{N}_{*}$ and bring to bear our calculations of the image of this map, from Proposition 3.1. All we need to do is to determine, in our set of generators for $\Omega^{Spin^{c}}_{*}/(2,\beta)$ through degree $33$, a maximal set of linear combinations of products of generators which generate $\beta$-torsion elements in $\Omega^{Spin^{c}}_{*}/(2)$, i.e., copies of $H\mathbb{F}_{2}$ rather than $ku_{(2)}$ in the Anderson–Brown–Peterson splitting.
@@
Re-running our inductive calculation from Proposition 3.1, but with this initial list for $D$, yields a set of $A$-module generators for $H^{*}(MSpin^{c};\mathbb{F}_{2})$ modulo $(Q_{0},Q_{1})$-torsion. Comparison of the lists produced by the first calculation and the second calculation, then using the translation matrix to translate back from the basis of Stiefel–Whitney monomials to the dual basis of partitions (i.e., Thom’s basis for $\mathfrak{N}_{*}$), gives us a set of generators for $\operatorname{{\rm im\>}}\left(\Omega^{Spin^{c}}_{*}\rightarrow\mathfrak{N}_{*}\right)$ in degrees $0,1,\dots,n$, and tells us, for each generator, whether it corresponds to a copy of $ku_{(2)}$ or of $H\mathbb{F}_{2}$ under the Anderson–Brown–Peterson splitting.
In degrees $\leq 33$, we find that an element $Y_{\lambda}\in\mathfrak{N}_{*}$ which is in the image of the map $\Omega^{Spin^{c}}_{*}\rightarrow\mathfrak{N}_{*}$ is $2$-torsion as long as the partition $\lambda$ includes an odd number. As described in section 1.1, the $\beta$-torsion elements of $\Omega^{Spin^{c}}_{*}$ are exactly the $2$-torsion elements. This yields the presentation for $\Omega^{Spin^{c}}_{*}$ in degrees $\leq 33$ in the statement of the theorem. ∎
@@
Manifold representatives for elements $Y_{n}$ in Thom’s partition basis for $\mathfrak{N}_{*}$ are as follows:
@@
Routine calculation using Stiefel–Whitney numbers. ∎
In [14], Milnor investigates whether every spin manifold is unorientedly cobordant to the square of an orientable manifold. He shows it is true for spin manifolds of dimension $\leq 23$. The ambiguity in dimension $24$ stems from the existence of an orientable manifold whose only nonzero Stiefel–Whitney numbers are $w_{4}w_{6}w_{7}^{2}$, $w_{6}^{4}$, $w_{4}^{6}$, $w_{4}^{3}w_{6}^{2}$, and $w_{4}^{2}w_{8}^{2}$. Milnor then poses the problem of whether a spin manifold of dimension $24$ exists with these nonzero Stiefel–Whitney numbers. Anderson–Brown–Peterson stated two years later [3] that, as a corollary of their main theorem, the lowest dimension in which there exists an element of $\text{Im}(\Omega^{Spin}_{*}\rightarrow\mathfrak{N}_{*})$ which is not the square of an orientable manifold is $24$ [3]. In Proposition 3.2 we calculated $\Omega^{Spin}_{*}$ in degrees through $31$ in terms of Thom’s partition basis, and in proposition 3.5 we have manifold representatives for ring-theoretic generators which suffice to generate everything in $\Omega^{Spin}_{24}$. Solving for an element of $\Omega^{Spin}_{24}$ with Milnor’s prescribed Stiefel–Whitney numbers, we find an explicit manifold of the kind Milnor asked for:
@@
The cobordism class $T_{24}+Y_{12}^{2}+Y_{10}^{2}Y_{2}^{2}+Y_{8}^{2}Y_{4}^{2}+Y_{8}^{2}Y_{2}^{4}+Y_{6}^{2}Y_{4}^{2}Y_{2}^{2}+Y_{5}^{4}Y_{2}^{2}+Y_{4}^{6}\in\text{Im}(\Omega^{Spin}_{*}\rightarrow\mathfrak{N}_{*})$ has nonzero Stiefel–Whitney numbers $w_{4}w_{6}w_{7}^{2}$, $w_{6}^{4}$, $w_{4}^{6}$, $w_{4}^{3}w_{6}^{2}$, and $w_{4}^{2}w_{8}^{2}$, and is represented by the manifold:
@@
In Theorem 3.4, we showed that, in degrees $\leq 33$, the cobordism classes $Y_{(5,5)},Y_{(9,9)},Y_{(11,11)},$ and $Y_{(13,13)}\in\mathfrak{N}_{*}$ lift to indecomposable $2$-torsion elements $Z_{10},Z_{18},Z_{22},$ and $Z_{26}$ in $\Omega^{Spin^{c}}_{*}$. The elements $Y_{(5,5)},Y_{(9,9)},\dots$ are precisely those of the form $Y_{i}^{2}\in\mathfrak{N}_{*}$ with $i$ odd and non-dyadic. It is natural to ask whether this pattern extends above degree $33$ as well. In Proposition 4.2, we show that something of this kind is true, if we use the Dold manifold $D_{i}$, from [10], rather than the cobordism class $Y_{i}$ from Thom’s partition basis for $\mathfrak{N}_{*}$. In low degrees, the algebraic relation between the Dold manifolds and the Thom generators for $\mathfrak{N}_{*}$ is as follows:
@@
The squares of odd-dimensional Dold manifolds represent the following polynomials in Thom’s generators $Y_{2},Y_{4},Y_{5},Y_{6},Y_{8},\dots$ for $\mathfrak{N}_{*}$:
@@
Dold [10] proves that there exists a minimal set of generators for the cobordism ring $\mathfrak{N}_{*}$ whose odd-degree elements are $D_{i}=P(2^{r}-1,s2^{r})$, where $i+1=2^{r}(2s+1)$ and $i$ is odd. Milnor proved that the map $\Omega^{U}_{*}\longrightarrow\mathfrak{N}_{*}$ maps onto all squares of elements in $\mathfrak{N}_{*}$. This map factors through $\Omega^{Spin^{c}}_{*}$, so $D_{i}\times D_{i}$ lifts to $\Omega^{Spin^{c}}_{2i}$ for all non-dyadic $i$. The mod $2$ cohomology of the Dold manifold $P(m,n)$ is given as a graded ring by
@@
with $c\in H^{1}(P(m,n);\mathbb{F}_{2})$ and $d\in H^{2}(P(m,n);\mathbb{F}_{2})$. The total Stiefel–Whitney class of $P(m,n)$ is
@@
To show that $D_{i}\times D_{i}$ lifts to the spin cobordism ring, one can carry out an algebraic calculation to show that none of the nonzero Stiefel–Whitney numbers of $D_{i}\times D_{i}$ are divisible by $w_{2}$. This is not a difficult calculation, but it is simpler to invoke the main result of Anderson’s paper [4]: the square of any orientable compact manifold is unorientedly cobordant to a spin manifold. Since $w_{1}$ vanishes on $D_{i}$, its square $D_{i}\times D_{i}$ must be in the image of the map $\Omega^{Spin}_{2i}\rightarrow\mathfrak{N}_{2i}$.
It follows easily from the structure of $ko_{*}$, and the Anderson–Brown–Peterson splitting of MSpin into a wedge of suspensions of $ko$, $ko\langle 2\rangle$, and $H\mathbb{F}_{2}$, that all elements of $\Omega^{Spin}_{*}$ in degrees $\equiv 2\mod 4$ are $2$-torsion. Hence, for odd $i$, any spin-cobordism class that lifts $D_{i}\times D_{i}\in\mathfrak{N}_{2i}$ must be $2$-torsion.
@@
To see that $D_{i}^{2}\in\Omega^{O}_{2i}$ lifts to an indecomposable element of $\Omega^{Spin}_{*}$, it is enough to observe that $\Omega^{Spin}_{*}/(2,\eta,\alpha,\beta)$ embeds into $\mathfrak{N}_{*}$, and since $D_{i}$ has second Stiefel–Whitney class $d\neq 0$, $D_{i}$ does not lift to $\Omega^{Spin}_{*}$. Hence the unique lift of $D_{i}^{2}$ to $\Omega^{Spin}_{*}/(2,\eta,\alpha,\beta)$ is indecomposable, hence any lift of $D_{i}^{2}$ to $\Omega^{Spin}_{*}$ is indecomposable. A completely analogous argument establishes that any lift of $D_{i}^{2}$ to $\Omega^{Spin^{c}}_{*}$ is also indecomposable. ∎
Since Dold [10] showed that $D_{i}\in\mathfrak{N}_{i}$ can be written as Thom’s generator $Y_{i}$ plus decomposables in the same degree, Proposition 4.2 tells us that some of the patterns exhibited in table (2) are not limited to degrees $\leq 33$, and indeed extend to all degrees. Namely, for odd non-dyadic $i$, if we write $Z_{2i}$ for a lift of $D_{i}\times D_{i}\in\mathfrak{N}_{2i}$ to an indecomposable $2$-torsion element of $\Omega^{Spin^{c}}_{2i}$ (guaranteed to exist by Proposition 4.2), and for even $i$ we write $Z_{2i}$ for a lift of $Y_{i}^{2}\in\mathfrak{N}_{2i}$ to an element of $\Omega^{Spin^{c}}_{2i}$, then we have:
@@
In this section, we show that the torsion-free generators $Z_{4},Z_{8},Z_{12},Z_{16},\dots$, together with the large nonunital subring of the $2$-torsion in $\Omega^{Spin^{c}}_{*}$, constructed in Theorem 4.3, suffice to generate a subring of the mod $2$ spinc-cobordism ring $\mathbb{F}_{2}\otimes_{\mathbb{Z}}\Omega^{Spin^{c}}_{*}$ which, although smaller than $\mathbb{F}_{2}\otimes_{\mathbb{Z}}\Omega^{Spin^{c}}_{*}$, is nevertheless uniformly $F$-isomorphic to all of $\mathbb{F}_{2}\otimes_{\mathbb{Z}}\Omega^{Spin^{c}}_{*}$. See Definition 1.1 for the definition of a uniform $F$-isomorphism.
@@
with $\beta$ the Bott element in degree $2$, with $y_{4i}$ in degree $4i$, and with $Z_{4j-2}$ in degree $24j-$ 2.
@@
Write $T$ for the ideal of $\Omega^{Spin^{c}}_{*}$ consisting of the $2$-torsion elements. Stong [23, Proposition 14] proved that $\left(\Omega^{Spin^{c}}_{*}/T\right)\otimes_{\mathbb{Z}}\mathbb{F}_{2}$ is a polynomial $\mathbb{F}_{2}$-algebra on generators $y_{2}$ and $y_{4},y_{8},y_{12},\dots$. Since $\Omega^{Spin^{c}}_{2}\cong\mathbb{Z}$ generated by $\beta=[\mathbb{C}P^{1}]$, Stong’s generator $y_{2}$ agrees modulo $2$ with the Bott element $\beta$. Hence $\left(\Omega^{Spin^{c}}_{*}/T\right)\otimes_{\mathbb{Z}}\mathbb{F}_{2}$ is isomorphic as a graded $\mathbb{F}_{2}$-algebra to $\mathbb{F}_{2}\left[\beta,y_{4i}:i\geq 1\right]$.

### v2

@@ 3. The $Spin^{c}$ bordism ring in low degrees
The image of the map $\Omega^{Spin}_{*}\longrightarrow\Omega^{O}_{*}$ agrees, in degrees $\leq 31$, with the subring of $MO_{*}$ generated by the elements
@@
While the individual element $Y_{i}\in\Omega^{O}_{i}$ does not generally lift to $\Omega^{Spin^{c}}_{i}$, its square $Y_{i}^{2}\in\Omega^{O}_{2i}$ does lift to the element $Z_{2i}\in\Omega^{Spin^{c}}_{2i}$. Consequently the squares of each of the monomials in each of the elements $T_{i}$ lift to $\Omega^{Spin^{c}}_{*}$. For $i=24,29,31,32$, and $33$, let $U_{2i}$ denote the element of $\Omega^{Spin^{c}}_{*}$ obtained by taking the definition of $T_{i}$ in Proposition 3.1 and replacing each instance of $Y_{n}$ with $Z_{2n}$. For example, (8) yields that
@@
The ring $\left(\Omega^{Spin^{c}}_{*}/T\right)\otimes_{\mathbb{Z}}\mathbb{F}_{2}$ was calculated by Stong [24, Proposition 11]: it is a polynomial $\mathbb{F}_{2}[\beta]$-algebra on generators in degrees $4,8,12,16,\dots$. Since $\tilde{T}$ is an ideal in $\Omega^{Spin^{c}}_{*}$, to calculate the product in the ring $\Omega^{Spin^{c}}_{*}\otimes_{\mathbb{Z}}\mathbb{F}_{2}$, it suffices to calculate
@@
Both of these types of products land in $\tilde{T}$. Since $\tilde{T}$ maps injectively under the map $\Omega^{Spin^{c}}_{*}\rightarrow\Omega^{O}_{*}$, we can embed $\Omega^{Spin^{c}}_{*}/(2,\beta)$ into $\Omega^{O}_{*}$ and bring to bear our calculations of the image of this map, from Proposition 3.1. All we need to do is to determine, in our set of generators for $\Omega^{Spin^{c}}_{*}/(2,\beta)$ through degree $33$, a maximal set of linear combinations of products of generators which generate $\beta$-torsion elements in $\Omega^{Spin^{c}}_{*}/(2)$, i.e., copies of $H\mathbb{F}_{2}$ rather than $ku_{(2)}$ in the Anderson–Brown–Peterson splitting.
@@
Re-running our inductive calculation from Proposition 3.1, but with this initial list for $D$, yields a set of $A$-module generators for $H^{*}(MSpin^{c};\mathbb{F}_{2})$ modulo $(Q_{0},Q_{1})$-torsion. Comparison of the lists produced by the first calculation and the second calculation, then using the translation matrix to translate back from the basis of Stiefel–Whitney monomials to the dual basis of partitions (i.e., Thom’s basis for $\Omega^{O}_{*}$), gives us a set of generators for $\operatorname{{\rm im\>}}\left(\Omega^{Spin^{c}}_{*}\rightarrow\Omega^{O}_{*}\right)$ in degrees $0,1,\dots,n$, and tells us, for each generator, whether it corresponds to a copy of $ku_{(2)}$ or of $H\mathbb{F}_{2}$ under the Anderson–Brown–Peterson splitting.
In degrees $\leq 33$, we find that an element $Y_{\lambda}\in\Omega^{O}_{*}$ which is in the image of the map $\Omega^{Spin^{c}}_{*}\rightarrow\Omega^{O}_{*}$ is $2$-torsion as long as the partition $\lambda$ includes an odd number. As described in section 1.1, the $\beta$-torsion elements of $\Omega^{Spin^{c}}_{*}$ are exactly the $2$-torsion elements. This yields the presentation for $\Omega^{Spin^{c}}_{*}$ in degrees $\leq 33$ in the statement of the theorem. ∎
Remark 3.5.
We marked one step in the proof of Theorem 3.4 with a footnote. In principle, that step in the calculation could go wrong, failing to identify all the $(Q_{0},Q_{1})$-torsion in $H^{*}(MSpin^{c};\mathbb{F}_{2})$, as follows: suppose there is some $\mathbb{F}_{2}$-linear combination of Stiefel–Whitney monomials in $H^{*}(MSpin^{c};\mathbb{F}_{2})$ which is $(Q_{0},Q_{1})$-torsion, but none of its summands are $(Q_{0},Q_{1})$-torsion. If this occurs, it would not be noticed by the method we describe, since our method only checks the $(Q_{0},Q_{1})$-torsion status of Stiefel–Whitney monomials.
We handle this by a very simple idea: we make the calculation as described, and after making the full calculation, we “check our answer” by comparing to the known additive structure of $\Omega^{Spin^{c}}_{*}$, as follows. After running our method, we compare the rank of our calculated $(Q_{0},Q_{1})$-torsion in each degree to the expected rank, using the known Poincaré series for the $2$-torsion in $\pi_{*}(MSpin^{c})$. If our method has failed to notice a linear combination of Stiefel–Whitney monomials which was $(Q_{0},Q_{1})$-torsion despite its summands not being $(Q_{0},Q_{1})$-torsion, then the rank of the $(Q_{0},Q_{1})$-torsion from our calculation will be too small.
We have never observed this mismatched rank to happen, i.e., it does not happen through degree $33$ in $MSpin^{c}_{*}$. If the rank mismatch were to ever occur, the fix is conceptually trivial, but computationally very hard: instead of populating $D$ with all the homogeneous $(Q_{0},Q_{1})$-torsion Stiefel–Whitney monomials at the start of the torsion calculation, we simply populate $D$ with the all the $(Q_{0},Q_{1})$-torsion Stiefel–Whitney polynomials at the start of the calculation, then re-run the calculation. This of course cannot miss any $(Q_{0},Q_{1})$-torsion in $H^{*}(MSpin^{c};\mathbb{F}_{2})$! Its disadvantage is simply that it is extremely computationally expensive, since the total number of homogeneous Stiefel–Whitney polynomials (not just monomials) in degrees $\leq d$ in $H^{*}(MSpin^{c};\mathbb{F}_{2})$ grows extremely quickly as $d$ grows. We carry out the calculations in the way we describe—i.e., initially populating $D$ by only the $(Q_{0},Q_{1})$-torsion Stiefel–Whitney monomials, rather than polynomials—to dramatically speed up calculation, and because we are able to check that the resulting answer in the end agrees with the answer we would have gotten with the much slower calculation using all the homogeneous Stiefel–Whitney polynomials.
@@
Manifold representatives for elements $Y_{n}$ in Thom’s partition basis for $\Omega^{O}_{*}$ are as follows:
@@
Routine calculation using Stiefel–Whitney numbers, since the unoriented cobordism class of a manifold is determined completely by its list of Stiefel–Whitney numbers. For the total Stiefel–Whitney classes of real projective spaces and of Dold manifolds, see [16, section 4 example 4] and [10, Satz 2], respectively. Once one has the total Stiefel–Whitney class of an $n$-dimensional manifold, one pairs with the top class to get the Stiefel–Whitney numbers, one such number for each homogeneous polynomial of degree $n$ in the Stiefel–Whitney classes $w_{1},\dots,w_{n}$. ∎
In [15], Milnor investigates whether every spin manifold is unorientedly cobordant to the square of an orientable manifold. He shows it is true for spin manifolds of dimension $\leq 23$. The ambiguity in dimension $24$ stems from the existence of an orientable manifold whose only nonzero Stiefel–Whitney numbers are $w_{4}w_{6}w_{7}^{2}$, $w_{6}^{4}$, $w_{4}^{6}$, $w_{4}^{3}w_{6}^{2}$, and $w_{4}^{2}w_{8}^{2}$. Milnor then poses the problem of whether a spin manifold of dimension $24$ exists with these nonzero Stiefel–Whitney numbers. Anderson–Brown–Peterson stated two years later [3] that, as a corollary of their main theorem, the lowest dimension in which there exists an element of $\text{Im}(\Omega^{Spin}_{*}\rightarrow\Omega^{O}_{*})$ which is not the square of an orientable manifold is $24$ [3]. In Proposition 3.2 we calculated $\Omega^{Spin}_{*}$ in degrees through $31$ in terms of Thom’s partition basis, and in proposition 3.6 we have manifold representatives for ring-theoretic generators which suffice to generate everything in $\Omega^{Spin}_{24}$. Solving for an element of $\Omega^{Spin}_{24}$ with Milnor’s prescribed Stiefel–Whitney numbers, we find an explicit manifold of the kind Milnor asked for:
@@
The cobordism class $T_{24}+Y_{12}^{2}+Y_{10}^{2}Y_{2}^{2}+Y_{8}^{2}Y_{4}^{2}+Y_{8}^{2}Y_{2}^{4}+Y_{6}^{2}Y_{4}^{2}Y_{2}^{2}+Y_{5}^{4}Y_{2}^{2}+Y_{4}^{6}\in\text{Im}(\Omega^{Spin}_{*}\rightarrow\Omega^{O}_{*})$ has nonzero Stiefel–Whitney numbers $w_{4}w_{6}w_{7}^{2}$, $w_{6}^{4}$, $w_{4}^{6}$, $w_{4}^{3}w_{6}^{2}$, and $w_{4}^{2}w_{8}^{2}$, and is represented by the manifold:
@@
In Theorem 3.4, we showed that, in degrees $\leq 33$, the cobordism classes $Y_{(5,5)},Y_{(9,9)},Y_{(11,11)},$ and $Y_{(13,13)}\in\Omega^{O}_{*}$ lift to indecomposable $2$-torsion elements $Z_{10},Z_{18},Z_{22},$ and $Z_{26}$ in $\Omega^{Spin^{c}}_{*}$. The elements $Y_{(5,5)},Y_{(9,9)},\dots$ are precisely those of the form $Y_{i}^{2}\in\Omega^{O}_{*}$ with $i$ odd and non-dyadic. It is natural to ask whether this pattern extends above degree $33$ as well. In Proposition 4.2, we show that something of this kind is true, if we use the Dold manifold $D_{i}$, from [10], rather than the cobordism class $Y_{i}$ from Thom’s partition basis for $\Omega^{O}_{*}$. In low degrees, the algebraic relation between the Dold manifolds and the Thom generators for $\Omega^{O}_{*}$ is as follows:
@@
The squares of odd-dimensional Dold manifolds represent the following polynomials in Thom’s generators $Y_{2},Y_{4},Y_{5},Y_{6},Y_{8},\dots$ for $\Omega^{O}_{*}$:
@@
Dold [10] proves that there exists a minimal set of generators for the cobordism ring $\Omega^{O}_{*}$ whose odd-degree elements are $D_{i}=P(2^{r}-1,s2^{r})$, where $i+1=2^{r}(2s+1)$ and $i$ is odd. In [15, Theorem 1], Milnor proved that the map $\Omega^{U}_{*}\longrightarrow\Omega^{O}_{*}$ maps onto all squares of elements in $\Omega^{O}_{*}$. This map factors through $\Omega^{Spin^{c}}_{*}$, so $D_{i}\times D_{i}$ lifts to $\Omega^{Spin^{c}}_{2i}$ for all non-dyadic $i$. By [10, Satz 1], the mod $2$ cohomology of the Dold manifold $P(m,n)$ is given as a graded ring by
@@
with $c\in H^{1}(P(m,n);\mathbb{F}_{2})$ and $d\in H^{2}(P(m,n);\mathbb{F}_{2})$. By [10, Satz 2], the total Stiefel–Whitney class of $P(m,n)$ is
@@
To show that $D_{i}\times D_{i}$ lifts to the spin cobordism ring, one can carry out an algebraic calculation to show that none of the nonzero Stiefel–Whitney numbers of $D_{i}\times D_{i}$ are divisible by $w_{2}$. This is not a difficult calculation, but it is simpler to invoke the main result of Anderson’s paper [4]: the square of any orientable compact manifold is unorientedly cobordant to a spin manifold. Since $w_{1}$ vanishes on $D_{i}$, its square $D_{i}\times D_{i}$ must be in the image of the map $\Omega^{Spin}_{2i}\rightarrow\Omega^{O}_{2i}$.
It follows easily from the structure of $ko_{*}$, and the Anderson–Brown–Peterson splitting of MSpin into a wedge of suspensions of $ko$, $ko\langle 2\rangle$, and $H\mathbb{F}_{2}$, that all elements of $\Omega^{Spin}_{*}$ in degrees $\equiv 2\mod 4$ are $2$-torsion. Hence, for odd $i$, any spin-cobordism class that lifts $D_{i}\times D_{i}\in\Omega^{O}_{2i}$ must be $2$-torsion.
@@
To see that $D_{i}^{2}\in\Omega^{O}_{2i}$ lifts to an indecomposable element of $\Omega^{Spin}_{*}$, it is enough to observe that $\Omega^{Spin}_{*}/(2,\eta,\alpha,\beta)$ embeds into $\Omega^{O}_{*}$, and since $D_{i}$ has second Stiefel–Whitney class $d\neq 0$, $D_{i}$ does not lift to $\Omega^{Spin}_{*}$. Hence the unique lift of $D_{i}^{2}$ to $\Omega^{Spin}_{*}/(2,\eta,\alpha,\beta)$ is indecomposable, hence any lift of $D_{i}^{2}$ to $\Omega^{Spin}_{*}$ is indecomposable. A completely analogous argument establishes that any lift of $D_{i}^{2}$ to $\Omega^{Spin^{c}}_{*}$ is also indecomposable. ∎
Since Dold [10] showed that $D_{i}\in\Omega^{O}_{i}$ can be written as Thom’s generator $Y_{i}$ plus decomposables in the same degree, Proposition 4.2 tells us that some of the patterns exhibited in table (2) are not limited to degrees $\leq 33$, and indeed extend to all degrees. Namely, for odd non-dyadic $i$, if we write $Z_{2i}$ for a lift of $D_{i}\times D_{i}\in\Omega^{O}_{2i}$ to an indecomposable $2$-torsion element of $\Omega^{Spin^{c}}_{2i}$ (guaranteed to exist by Proposition 4.2), and for even $i$ we write $Z_{2i}$ for a lift of $Y_{i}^{2}\in\Omega^{O}_{2i}$ to an element of $\Omega^{Spin^{c}}_{2i}$, then we have:
@@
There is a subring of the mod $2$ spinc-cobordism ring $\Omega^{Spin^{c}}_{*}\otimes_{\mathbb{Z}}\mathbb{F}_{2}$ which is generated by
  • the large nonunital subring of the $2$-torsion in $\Omega^{Spin^{c}}_{*}$ constructed in Theorem 4.3,
  • and Stong’s generators $y_{4},y_{8},y_{12},y_{16},\dots$ of the mod-torsion mod-$2$ spinc-cobordism ring $(\Omega^{Spin^{c}}_{*}/tors)\otimes_{\mathbb{Z}}\mathbb{F}_{2}$.
This subring of $\Omega^{Spin^{c}}_{*}\otimes_{\mathbb{Z}}\mathbb{F}_{2}$ is strictly smaller than $\Omega^{Spin^{c}}_{*}\otimes_{\mathbb{Z}}\mathbb{F}_{2}$ itself. However, we will now show that this subring is uniformly $F$-isomorphic to $\Omega^{Spin^{c}}_{*}\otimes_{\mathbb{Z}}\mathbb{F}_{2}$. See Definition 1.2 for the definition of a uniform $F$-isomorphism.
@@
with $\beta$ the Bott element in degree $2$, with $y_{4i}$ in degree $4i$, and with $Z_{4j-2}$ in degree 4j24j-2.
@@
Write torstors for the ideal of $\Omega^{Spin^{c}}_{*}$ consisting of the $2$-torsion elements. Stong [24, Proposition 14] proved that $\left(\Omega^{Spin^{c}}_{*}/tors\right)\otimes_{\mathbb{Z}}\mathbb{F}_{2}$ is a polynomial $\mathbb{F}_{2}$-algebra on generators $y_{2}$ and $y_{4},y_{8},y_{12},\dots$. Since $\Omega^{Spin^{c}}_{2}\cong\mathbb{Z}$ generated by $\beta=[\mathbb{C}P^{1}]$, Stong’s generator $y_{2}$ agrees modulo $2$ with the Bott element $\beta$. Hence $\left(\Omega^{Spin^{c}}_{*}/tors\right)\otimes_{\mathbb{Z}}\mathbb{F}_{2}$ is isomorphic as a graded $\mathbb{F}_{2}$-algebra to $\mathbb{F}_{2}\left[\beta,y_{4i}:i\geq 1\right]$.